OpenStudy (anonymous):

Consider the function f(x)=3x-8. I found the inverse. I got x=x+8/3. what is the composition that show f(x) and g(x) are inverses of eachother. Please help. I'm desperate

OpenStudy (loser66):

First off, the inverse of f(x) =3x +8 is \(f^{-}(x) = \dfrac{x+8}{3}\) not x = x+8/3 as you said. Got it?

OpenStudy (anonymous):

That is the same thing I have.. F^1 x +8/3... But if your going to be mean you dont have to answer my question.

OpenStudy (loser66):

And if you name it as g(x) , then it is simple to show they are inverse of each other

OpenStudy (anonymous):

How....

OpenStudy (loser66):

Ok, I have to correct your mistake before doing something else. But if it frustrates you, I am sorry.

OpenStudy (loser66):

by showing f(g(x) = x.

OpenStudy (loser66):

and g(f(x) =x

OpenStudy (anonymous):

is that all I have to say...?

OpenStudy (loser66):

What do you expect? you say and prove what you say.

OpenStudy (anonymous):

I would appreciate it if you gave me an answer i wasn't unsure of...

OpenStudy (anonymous):

I am seeking help. Not judgement

OpenStudy (loser66):

f(x) = 3x -8 g(x) =\(\dfrac{x+8}{3}\) so that f(g(x) ) = 3 (g(x)) -8 = 3\(\dfrac{x+8}{3}\) -8 =x

OpenStudy (loser66):

do the same with g(f(x)). you will get g(f(x))= x then conclude f, g are inverse of each other. gtg. good luck